# WinkTPG: An Execution Framework for Multi-Agent Path Finding Using Temporal Reasoning

### v1

@@ Abstract
Planning collision-free paths for a large group of agents is a challenging problem with numerous real-world applications. While recent advances in Multi-Agent Path Finding (MAPF) have shown promising progress, standard MAPF algorithms rely on simplified kinodynamic models, preventing agents from directly following the generated MAPF plan. To bridge this gap, we propose kinodynamic Temporal Plan Graph Planning (kTPG), a multi-agent speed optimization algorithm that efficiently refines a MAPF plan into a kinodynamically feasible plan while accounting for uncertainties and preserving collision-freeness. Building on kTPG, we propose Windowed kTPG (WinkTPG), a MAPF execution framework that incrementally refines MAPF plans using a window-based mechanism, dynamically incorporating agent information during execution to reduce uncertainty. Experiments show that WinkTPG can generate speed profiles for up to 1,000 agents in 1 second and improves solution quality by up to 51.7% over existing MAPF execution methods.
1 Introduction
@@
Coordinating agents without collisions is important for applications in scenarios like automated warehouses (Wurman, D’Andrea, and Mountz 2008), traffic intersections (Li et al. 2023), and airports (Morris et al. 2016). Multi-Agent Path Finding (MAPF) methods offer significant advantages for solving this problem, such as scalability and (bounded) optimality guarantees. Yet, applying MAPF in these domains remains challenging: most MAPF methods ignore real-world factors like kinodynamics (e.g., speed and acceleration limits) and uncertainty (e.g., execution delay), making it impractical for agents to strictly follow the planned paths.
@@
One promising solution is to postprocess the plan generated by MAPF methods. Hönig et al. [2016] introduce Temporal Plan Graph (TPG), which refines the speed profiles of agents to meet their speed limits while preserving the passing orders of agents at different locations. While TPG ensures collision-free and deadlock-free, it cannot handle kinodynamic constraints beyond speed limits. That is, it may generate speed profiles with infinitely large accelerations, making them infeasible for agents to follow. Two extensions of TPG overcome this issue: Kinodynamic Network (KDN) (Zhang et al. 2021) and Action Dependency Graph (ADG) (Hönig et al. 2019). KDN incorporates kinodynamic constraints into TPG but suffers from limited scalability and cannot handle uncertainty. In contrast, ADG is a distributed execution framework using TPG that lets each agent plan its speed profile. However, as we later show, ADG is conservative, frequently forcing agents to decelerate unnecessarily.
To overcome this limitation, we propose Windowed Kinodynamic Temporal Plan Graph planning (WinkTPG), an efficient execution framework for MAPF. Given a set of collision-free paths, we use a TPG to encode their passing orders at different locations. Instead of taking the conservative strategy used by ADG, WinkTPG uses our proposed kinodynamic planner, kTPG, to plan speed profiles that leverage the movement capabilities of agents to avoid unnecessary deceleration. Specifically, kTPG first transforms the passing order constraints into reserved time intervals, during which agents must traverse specific locations. Then, for each agent, kTPG generates a speed profile that satisfies both reserved intervals and kinodynamic constraints. To ensure the safety of those speed profiles with respect to execution time uncertainty, we add a safety margin mechanism to kTPG. Moreover, WinkTPG uses a windowed mechanism to enhance robustness and efficiency during execution. This mechanism uses updated information during execution to reduce uncertainty and improve overall performance through dynamic replanning. Theoretically, we prove that kTPG and WinkTPG are complete. Empirically, we evaluate them on the standard MAPF benchmark with two robot models and various uncertainty levels. WinkTPG shows up to 51.7% improvement in terms of solution quality compared to ADG. Moreover, WinkTPG generates speed profiles for 1,000 agents within 1 second, whereas KDN requires up to 300 seconds and scales only to 200 agents.
2 Preliminaries
@@
Before describing our approach we first define the MAPF problem and the temporal plan graph that encodes the MAPF plan, and review related works on executing MAPF plans.
Figure 1: Example of MAPF problem, plan, and a TPG.
2.1 MAPF Problem
@@
As shown in Fig. 1 (a), a MAPF problem (Stern et al. 2019) contains an undirected graph $\mathcal{G}_{U}=(\mathcal{V}_{U},\mathcal{E}_{U})$ and $I$ agents $\mathcal{A}=\{a_{1},\dots,a_{I}\}$, each with a start $q_{i}^{s}\in\mathcal{V}_{U}$ and goal $q_{i}^{g}\in\mathcal{V}_{U}$. Time is discretized, and at each timestep, an agent can either move to an adjacent vertex or remain at its current vertex. A MAPF plan is a set of collision-free paths $\mathcal{P}=\{p_{1},\dots,p_{I}\}$, where the path of $a_{i}$ is a sequence of time-annotated vertices: $p_{i}=\{(q_{i}^{0},t_{0}),\dots,(q_{i}^{z^{i}},t_{z^{i}})\}$. Here, $q_{i}^{0}=q_{i}^{s},q_{i}^{z^{i}}=q_{i}^{g},q_{i}^{k}\in\mathcal{V}_{U}$ and $(q_{i}^{{k-1}},q_{i}^{{k}})\in\mathcal{E}_{U}$ for all $k\in\{1,\dots,z^{i}\}$. $\mathcal{P}$ is defined as collision-free if it contains No vertex collisions - agents must not occupy the same vertex at the same time, and No edge collisions - agents must not swap vertices simultaneously. To avoid confusion with the TPG, we refer to vertices in $\mathcal{V}_{U}$ as locations.
2.2 Temporal Plan Graph (TPG)
@@
A Temporal Plan Graph (TPG), as shown in Fig. 1 (b), is a directed acyclic graph $\mathcal{G}=(\mathcal{V},\mathcal{E}_{1},\mathcal{E}_{2})$ that represents a MAPF plan by capturing precedence relationships for visiting locations. The vertex set $\mathcal{V}=\{v_{i}^{k}:k\in[0,z^{i}],a_{i}\in\mathcal{A}\}$ comprises all locations to be visited sequentially by agent $a_{i}$ (i.e., $loc(v_{i}^{k})\in\mathcal{V}_{U}$). The edge set $\mathcal{E}_{1}$ includes Type-1 edges, which capture sequential precedence within the path of an agent. A Type-1 edge $(v_{i}^{k},v_{i}^{k+1})$ ensures that $a_{i}$ visits $v_{i}^{k}$ before $v_{i}^{k+1}$. The edge set $\mathcal{E}_{2}$ consists of Type-2 edges, which enforce the order in which agents visit shared locations. For any two vertices $v_{j}^{s}$ and $v_{i}^{k}$ such that $loc(v_{j}^{s})=loc(v_{i}^{k})$, a Type-2 edge $(v_{j}^{s+1},v_{i}^{k})$ ensures that $a_{i}$ can enter $v_{i}^{k}$ only after $a_{j}$ leaves $v_{j}^{s}$ and reaches $v_{j}^{s+1}$.
@@
Figure 2: (a) MAPF instance with a green agent (R1) moving from C2 to A2, and a blue agent (R2) moving from B5 to B1. Both agents have an initial speed of $U_{max}$. (b) TPG generated by the central server, where the yellow edge indicates that R2 can only move to B2 after R1 has left it. (c) ADG-generated and optimal speed profiles for R2.

### v2

@@ Abstract
Planning collision-free paths for a large group of agents is a challenging problem in many real-world applications. While recent advances in Multi-Agent Path Finding (MAPF) have shown promising progress, standard MAPF planners continue to rely on simplified kinodynamic models, preventing agents from directly following the generated MAPF plan. To bridge this gap, we propose kinodynamic Temporal Plan Graph planning (kTPG), a multi-agent speed optimization algorithm that efficiently refines a MAPF plan into a set of kinodynamically feasible speed profiles. We further incorporate execution timing uncertainty models and provide deterministic guarantees under bounded uncertainty models and probabilistic guarantees under stochastic models. Building on kTPG, we propose Windowed kTPG (WinkTPG), a MAPF execution framework that incrementally refines MAPF plans using a window-based mechanism, dynamically incorporating agent information during execution to reduce uncertainty. Experiments show that WinkTPG can generate speed profiles for up to 1,000 agents within 1 second and improves solution quality by up to 51.7% over existing MAPF execution methods. We further validate WinkTPG in high-fidelity physics-based simulation and on real-world robots.
Note to Practitioners
The motivation of this article originates from the need to execute large-scale multi-agent path-planning solutions reliably in practical applications such as warehouse logistics, industrial material transport, and factory or airport automation. Although discrete MAPF planners can generate conflict-free paths, their solutions are often not directly executable by real robots due to kinodynamic constraints and execution timing uncertainties. This article develops WinkTPG, a windowed execution framework that converts MAPF plans into safe and kinodynamically feasible speed profiles while maintaining precedence relations implied by the discrete plan. WinkTPG further employs receding-horizon replanning to correct timing deviations caused by disturbances or sensing noise, making it suitable for practitioners deploying multi-robot fleets in constrained, time-critical environments.
I Introduction
@@
Collision-free coordination of multiple robotic agents operating in the same physical space is an important task in numerous real-world applications, including automated warehouses [29, 19], traffic intersections [13], and airports [21]. Multi-Agent Path Finding (MAPF) methods offer significant advantages for solving this problem, such as scalability and (bounded) optimality guarantees. However, applying MAPF in these domains remains challenging: most MAPF methods ignore real-world factors like kinodynamics (e.g., speed and acceleration limits) and execution timing uncertainty (i.e., variability in traversal times due to controller noise, actuation latency, and slippage), making it impractical for agents to strictly follow the planned paths.
@@
One promising solution is to postprocess the plan generated by MAPF methods. Honig et al. [10] introduce the Temporal Plan Graph (TPG), which refines the speed profiles of agents to meet their speed limits while preserving the passing orders of agents at different locations. While TPG ensures collision-free and deadlock-free movement, it cannot handle kinodynamic constraints beyond speed limits. That is, it may generate speed profiles with infinitely large accelerations, making them infeasible for agents to follow. Two extensions of TPG overcome this problem: the Kinodynamic Network (KDN) [32] and the Action Dependency Graph (ADG) [9]. KDN incorporates kinodynamic constraints into TPG but suffers from limited scalability and cannot handle execution timing uncertainty. In contrast, ADG is a scalable execution framework that transforms a MAPF plan into a dependency graph and coordinates agents through event-triggered execution. However, each agent in an ADG plans its own speed profile independently, without accounting for the planned or predicted motion of other agents. Instead of jointly coordinating speed profiles, each agent reacts only after its dependencies are formally resolved, ignoring the future motion of other agents even when it could be anticipated. Consequently, ADG often under-utilizes available temporal flexibility, causing unnecessary deceleration and reducing overall execution efficiency even when globally coordinated motion would be feasible.
To overcome this limitation, we propose Windowed Kinodynamic Temporal Plan Graph planning (WinkTPG), an efficient execution framework for MAPF. Unlike event-driven methods such as ADG, which make execution decisions for each agent independently without global coordination, WinkTPG jointly plans speed profiles for all agents under the temporal structure induced by the TPG. Given a set of collision-free paths, WinkTPG first constructs a TPG to encode the passing orders of agents at shared locations. It then applies the proposed kinodynamic planner, kTPG, to generate coordinated speed profiles for all agents. The key idea of kTPG is to use reserved intervals to allocate shared locations to competing agents over time. These intervals specify when agents are allowed to traverse particular locations, based on kinodynamic reachability and the precedence structure of the TPG. By propagating and updating reserved intervals across agents, kTPG captures global temporal coupling among agents rather than making purely local, event-triggered decisions. This enables kTPG to optimize the motions of agents in a coordinated manner and to reduce the conservative behavior commonly incurred by ADG. To improve robustness under modeled execution timing uncertainty, we incorporate a safety-margin mechanism into kTPG. Under bounded execution timing uncertainty models, our method guarantees deterministic precedence satisfaction and collision-free execution; under a stochastic uncertainty model, it yields probabilistic guarantees.
To enhance robustness and efficiency during execution, WinkTPG couples kTPG with the use of a windowed execution mechanism. This mechanism uses updated information during execution to reduce uncertainty and improve overall performance through dynamic replanning. Theoretically, we prove that kTPG and WinkTPG are complete. Empirically, we evaluate them on the standard MAPF benchmarks with two robot models and three execution timing uncertainty models. WinkTPG shows up to 51.7% improvement in terms of solution quality compared to ADG. Moreover, WinkTPG generates speed profiles for 1,000 agents within 1 second, whereas KDN may take over 300 seconds to generate solutions for only 200 agents. Finally, we demonstrate the real-world feasibility of WinkTPG by validating it in a high-fidelity physics-based simulation environment and on real-world robots.
Fig. 1 provides an overview of the proposed framework and also serves as a guide to the organization of this paper. Starting from a discrete MAPF plan, we first formulate the multi-agent execution problem (MAEP) and introduce kTPG in Section III. Then, in Section IV we present the uncertainty-aware extension of MAEP and discuss how execution timing uncertainty is incorporated into the planning process. Building on this formulation, we next describe WinkTPG, a windowed execution strategy that enables online replanning using updated execution information in Section V. Theoretical and Empirical results are presented in Sections VI and VII, respectively.
II Preliminaries
@@
Before describing our method, we first define the MAPF problem and the temporal plan graph that encodes the MAPF plan, and review related works on executing MAPF plans.
Figure 1: Overview of WinkTPG. WinkTPG produces robust speed profiles for all agents given a MAPF plan. A discrete MAPF plan is first converted into a Temporal Planning Graph (TPG). Kinodynamic constraints are incorporated to form kTPG (Sec. III), and execution uncertainty is modeled in kTPGu (Sec. IV). During execution, WinkTPG (Sec. V) updates temporal windows based on real-time execution feedback, enabling adaptive and less conservative execution.
Figure 2: Example of a MAPF instance, a MAPF plan, and a TPG.
II-A MAPF Problem
@@
As shown in Fig. 2 (a), a MAPF problem [26] contains an undirected graph $\mathcal{G}_{U}=(\mathcal{V}_{U},\mathcal{E}_{U})$ and $I$ agents $\mathcal{A}=\{a_{1},\dots,a_{I}\}$, each with a start $q_{i}^{s}\in\mathcal{V}_{U}$ and goal $q_{i}^{g}\in\mathcal{V}_{U}$. Time is discretized, and at each timestep, an agent can either move to an adjacent vertex or remain at its current vertex. A MAPF plan is a set of collision-free paths $\mathcal{P}=\{p_{1},\dots,p_{I}\}$, where the path of $a_{i}$ is a sequence of time-annotated vertices: $p_{i}=\{(q_{i}^{0},t_{0}),\dots,(q_{i}^{z^{i}},t_{z^{i}})\}$. Here, $q_{i}^{0}=q_{i}^{s},q_{i}^{z^{i}}=q_{i}^{g}$, and for all $k\in\{1,\dots,z^{i}\},q_{i}^{k}\in\mathcal{V}_{U}$ and $(q_{i}^{{k-1}},q_{i}^{{k}})\in\mathcal{E}_{U}$. $\mathcal{P}$ is defined as collision-free if it contains no vertex collisions: agents must not occupy the same vertex at the same timestep, and no edge collisions: agents must not swap vertices at the same timestep. To avoid confusion with the TPG, we henceforth refer to vertices in $\mathcal{V}_{U}$ as locations. To be compatible with the TPG representation, we assume that the MAPF plan is also cycle-conflict-free (i.e., no subset of agents swap each other’s locations at the same timestep [26]).
II-B Temporal Plan Graph (TPG)
@@
A Temporal Plan Graph (TPG), as shown in Fig. 2 (b), is a directed acyclic graph $\mathcal{G}=(\mathcal{V},\mathcal{E}_{1},\mathcal{E}_{2})$ that represents a MAPF plan $\mathcal{P}$ by capturing precedence relationships for visiting locations. The vertex set $\mathcal{V}=\{v_{i}^{k}:k\in[0,z^{i}],a_{i}\in\mathcal{A}\}$ comprises all locations to be visited sequentially by agent $a_{i}$ (i.e., $loc(v_{i}^{k})=q_{i}^{k}$). The edge set $\mathcal{E}_{1}$ includes Type-1 edges, which capture sequential precedence within the path of an agent. A Type-1 edge $(v_{i}^{k},v_{i}^{k+1})$ ensures that $a_{i}$ visits $v_{i}^{k}$ before $v_{i}^{k+1}$. The edge set $\mathcal{E}_{2}$ consists of Type-2 edges, which enforce the order in which agents visit shared locations. For any two vertices $v_{j}^{s}$ and $v_{i}^{k}$ such that $loc(v_{j}^{s})=loc(v_{i}^{k})$, a Type-2 edge $(v_{j}^{s+1},v_{i}^{k})$ ensures that $a_{i}$ can enter $v_{i}^{k}$ only after $a_{j}$ leaves $v_{j}^{s}$ and reaches $v_{j}^{s+1}$.
II-C Related Works
MAPF methods compute collision-free paths for large groups of agents. Leading methods, such as PBS [18], PIBT [23], and EECBS [14], can find collision-free paths for thousands of agents. However, these methods assume all agents move synchronously, which is unrealistic in real-world scenarios due to their reliance on simplified, discretized kinodynamic models. This limitation is also reflected in prior studies that demonstrate the execution of MAPF plans on physical multi-robot platforms [4, 31, 25, 12, 28], highlighting the practical importance of bridging the gap between discrete MAPF plans and continuous real-world execution. To address this issue, some methods [2, 30, 20] integrate more accurate kinodynamic models during planning. However, these methods often face scalability challenges due to the increased computational complexity introduced by the kinodynamic models. Another group of methods [10, 32] first generates a MAPF plan with a simplified kinodynamic model and subsequently derives speed profiles with the accurate kinodynamic model based on these paths. For instance, KDN [32] employs a TPG to encode the precedence constraints defined in the MAPF plan. It then uses a Mixed-Integer Linear Programming (MILP) solver to find kinodynamically feasible speed profiles for agents based on the TPG.
To handle execution timing uncertainty, some methods [3, 24] account for this uncertainty during planning. These methods first build a model of the execution timing uncertainty and then plan collision-free paths, ensuring that agents remain collision-free even when time uncertainties occur as described by the models. While effective in some cases, these methods often lead to poor performance due to being overly conservative during the planning stage, as they cannot access the actual delays that occur during execution. Additionally, these methods are often limited in scalability.
Introducing an execution framework to MAPF is a promising method that enables scalable coordination while satisfying agent kinodynamic constraints and ensuring robust execution. Action Dependency Graph (ADG) [9] uses a TPG to ensure the passing order and defines an execution framework for speed profile generation. Specifically, it categorizes the vertices in the graph into three statuses: staged, enqueued, and finished. A vertex is marked as staged if the agent cannot move to it yet. It transitions to enqueued when the agent can safely move here. This happens only if it has no preceding vertices or if the following conditions are met: (1) its preceding vertex connected by a Type-1 edge is enqueued or finished, and (2) all preceding vertices connected by Type-2 edges are finished. Once a vertex transitions to enqueued, ADG generates a speed profile for the agent to reach the vertex. A vertex is marked as finished when the agent has reached it. ADG dynamically updates vertex statuses during execution and provides robustness guarantees under execution timing uncertainty.
Although ADG provides robustness guarantees, it does not explicitly account for the speed profiles and execution timing of other agents when reasoning dependencies. Because dependency satisfaction is determined solely by predecessor completion events, ADG cannot reason about when a shared location will become temporally feasible based on ongoing execution of other agents. As a result, it must adopt a conservative strategy in which an agent waits until all required predecessors have formally finished, even when their planned speed profiles already imply conflict-free traversal. This conservative strategy can lead to unnecessary deceleration. As shown in Fig. 3, R2 (blue) must wait until R1 (green) leaves cell B2 before it can proceed. Although it is evident early on that R1 will leave B2 before R2 arrives, ADG does not allow R2 to commit to the optimal speed profile until the dependency is officially satisfied. This results in R2 decelerating prematurely (top chart in (c)), compared to the optimal case (bottom chart) where it could have maintained maximum speed. The problem is worsened by potential communication delays in real-world settings, where agents rely on frequent updates from a central server. Consequently, this overly conservative strategy can result in poor performance in many scenarios.
@@ II-C Related Works
Figure 3: (a) MAPF instance with a green agent (R1) moving from C2 to A2, and a blue agent (R2) moving from B5 to B1. Both agents have an initial speed of $v_{max}$. (b) TPG generated by the central server, where the yellow edge indicates that R2 can only move to B2 after R1 has left it. (c) ADG-generated and optimal speed profiles for R2.
III Kinodynamic TPG (kTPG)
Section III-A below specifies the problem addressed in this section. In Sections IV and V, we will extend this definition to account for execution timing uncertainty during agent movement.
III-A Multi-Agent Execution Problem (MAEP)
Given a MAPF plan, the objective of MAEP [32] is to generate a set of speed profiles for all agents to follow during execution. These profiles must satisfy the kinodynamic constraints imposed by the agents’ physical models, respect the precedence relationships defined in the MAPF plan, and preserve collision-free movement. The objective is to minimize the total time for all agents to reach their goals.
Let $A=\{a_{1},\dots,a_{I}\}$ denote the set of agents defined in Section II-A, now operating in a continuous workspace $\mathcal{W}\subset\mathbb{R}^{d}$, with $d\in\{2,3\}$. Given a MAPF plan $\mathcal{P}=\{p_{1},\dots,p_{I}\}$, we reinterpret each discrete path $p_{i}$ as a sequence of waypoints in continuous space. Specifically, for agent $a_{i}$, its path is given by $p_{i}=\big\{(q_{i}^{0},t_{i}^{0}),(q_{i}^{1},t_{i}^{1}),\dots,(q_{i}^{z_{i}},t_{i}^{z_{i}})\big\}$, where each $q_{i}^{k}$ corresponds to a vertex in the MAPF graph $\mathcal{G}_{U}=(\mathcal{V}_{U},\mathcal{E}_{U})$ and represents a location in $\mathcal{W}$. During execution, each agent is required to spatially follow its assigned path, i.e., visit the waypoints in order and move along straight-line segments between consecutive waypoints. To execute its path in continuous space, each agent $a_{i}$ follows a continuous trajectory $x_{i}(\cdot)$ that must satisfy all kinodynamic and safety constraints imposed by its physical model. We capture these limits by an admissible trajectory set:
$\mathcal{X}_{i}=\Big\{x_{i}:[0,\infty)\to\mathbb{R}^{d_{i}}\;\Big|\;C_{i}\big(x_{i}(\cdot)\big)\leq 0\Big\},$
where $d_{i}$ is the dimension of the state space of agent $a_{i}$, and $C_{i}(\cdot)$ is a constraint operator encoding the kinodynamic feasibility requirements of $a_{i}$, including bounds on velocity, acceleration, jerk, actuator capabilities, and other safety constraints. The operator $C_{i}(\cdot)$ may represent pointwise-in-time constraints (e.g., $\|\dot{x}_{i}(t)\|\leq v_{\max}$, $\|\ddot{x}_{i}(t)\|\leq a_{\max}$) as well as integral or state-dependent conditions. A trajectory is considered kinodynamically feasible if and only if it belongs to $\mathcal{X}_{i}$. We assume that waiting at waypoints is always kinodynamically feasible under $C_{i}(\cdot)$. We assume that the spatial paths produced by the MAPF planner are executable, in the sense that for each agent $a_{i}$ there exists at least one continuous trajectory $x_{i}(\cdot)\in\mathcal{X}_{i}$ that follows the assigned path $p_{i}$.
Although agents move in continuous space, potential inter-agent collisions can only occur at a finite set of shared waypoints defined by the MAPF plan. We therefore use the underlying MAPF graph $\mathcal{G}_{U}$ to reason about collision avoidance during execution. This graph abstraction allows collision avoidance to be enforced through temporal constraints at waypoints, rather than continuous-space reasoning. We define the reach time $t_{i}^{r}(q_{i}^{k})$ as the time at which $a_{i}$ arrives at waypoint $q_{i}^{k}$, and the corresponding leave time $t_{i}^{l}(q_{i}^{k})=t_{i}^{r}(q_{i}^{k+1})$ as the time at which it departs toward the next waypoint. We say $a_{i}$ occupies waypoint $q_{i}^{k}$ during $t_{i}^{r}(q_{i}^{k}),t_{i}^{l}(q_{i}^{k})]$.
We define a trajectory set $\{x_{i}(t)\}_{i=1}^{I}$ as valid for $\mathcal{P}$ if it satisfies:
  1. 1.
    Precedence Constraints: For every pair of consecutive waypoints assigned to an agent, the execution must respect the ordering implied by the MAPF plan:
    $t^{r}_{i}(q_{i}^{k})<t^{r}_{i}(q_{i}^{k+1}),\qquad\forall i,\forall k,$
    If agents $a_{i}$ and $a_{j}$ both visit $q$, and $a_{i}$ is assigned to visit $q$ before $a_{j}$, then
    $t^{l}_{i}(q)\leq t^{r}_{j}(q).$
  2. 2.
    Kinodynamic Feasibility:
    $x_{i}(\cdot)\in\mathcal{X}_{i},\qquad\forall i,\ \forall t.$
Note that the precedence constraint is sufficient to prevent both vertex and edge collisions. Specifically, the constraint $t^{l}_{i}(q)\leq t^{r}_{j}(q)$ ensures that the occupancy intervals of agents at any shared location do not overlap, thereby preventing vertex collisions. Furthermore, an edge collision would require two agents to traverse the same undirected edge in opposite directions over overlapping time intervals. Such a situation necessarily implies overlapping occupancy of at least one shared endpoint. Since the MAPF plan is collision-free, and since the execution preserves waypoint order while prohibiting overlapping occupancy at any vertex, opposite-direction traversal cannot occur.
Objective. Given the MAPF plan $\mathcal{P}$ and the kinodynamic models $\{\mathcal{X}_{i}\}$, the Multi-Agent Execution Problem is
$\min_{\{x_{i}(\cdot)\}}$ $\sum_{i=1}^{I}t^{r}_{i}(q_{i}^{z_{i}})$
s.t. $\{x_{i}(\cdot)\}\text{ is valid for }\mathcal{P}.$
III-B kTPG Overview
The high-level idea of kTPG is to encode the precedence constraints in a MAPF plan into a temporal plan graph (TPG) $\mathcal{G}$ and associate each vertex with a reserved interval. Beyond encoding precedence relations, kTPG synthesizes execution timing information from the speed profiles of multiple agents together with kinodynamic reachability constraints into these reserved intervals. Each reserved interval therefore specifies when an agent can arrive at a location without conflicting with others or violating its kinodynamic constraints. kTPG generates speed profiles for individual agents independently, ensuring that each agent visits the vertices along its path within the corresponding reserved intervals.
All reserved intervals are initialized to $0,\infty)$, and kTPG iteratively refines both the intervals and the associated speed profiles. When a speed profile is fixed for one agent, the induced occupancy timing constrains the feasible arrival windows of dependent vertices, and these constraints are propagated through $\mathcal{G}$. The refinement process continues until the reserved intervals of vertices corresponding to the same location do not overlap, ensuring collision-free execution. We say that a set of speed profiles satisfies the edges in $\mathcal{G}$ if the occupancy intervals induced by the speed profiles do not violate the precedence constraints encoded by the edges. Satisfaction of all edges in the TPG guarantees that the precedence relations defined in the MAPF plan are preserved.
Algorithm 1 summarizes the overall procedure of kTPG. Using the example in Fig. 4, we illustrate a single iteration of this process. Each iteration begins by selecting an agent to refine its speed profile, following the selection rule in Section III-C; in this case, $a_{2}$ is selected. We define a Type-2 edge as satisfied if the reserved intervals at its associated vertices do not overlap and the visiting precedence is preserved. Otherwise, it is conflicting. All Type-2 edges in our example are conflicting in this iteration. We then generate a speed profile for $a_{2}$ from its first to last unlocked vertex, i.e., those vertices without any incoming conflicting edges, using the method in Section III-D. These vertices are shown as dark green circles in Fig. 4 (b). Next, we identify the conflicting edges originating from these unlocked vertices (i.e., the red edge in Fig. 4 (b)). We use the generated speed profile to derive a split time for the target vertex (here, C2) and update the reserved intervals for both agents accordingly so that the intervals no longer overlap (see Section III-E). This update allows us to satisfy a conflicting edge and unlock additional vertices of other agents, as shown in Fig. 4 (c). In the next iteration, we use the newly unlocked vertices to satisfy more Type-2 edges by selecting another agent and updating its speed profile. This process repeats until all Type-2 edges are satisfied, at which point we have complete, collision-free speed profiles for all agents.